OpenStudy (iesmee):

find solution (x,y,z) 3x+5-3z=27 -3x-y+2z=-3 -2x+2y+6z=14

OpenStudy (iesmee):

@satellite73

OpenStudy (iesmee):

@wio

OpenStudy (iesmee):

@dan815

OpenStudy (amoodarya):

did you mean ? 3x+5y-3z=27 -3x-y+2z=-3 -2x+2y+6z=14

OpenStudy (amoodarya):

find x from the 1st equation and out into 2nd and 3rd then find y from one of them and put into other to find z now solve for z then go backward to find y then for x

OpenStudy (anonymous):

is this a simultaneous equation

OpenStudy (anonymous):

Gauss-Jordan Matrix Elimination.

OpenStudy (amoodarya):

-2x+2y+6z=14 /2 -x+y+3z=7 y+3z-7=x now put x=y+3z-7 in the 1st and 2nd equation simplify solve for y,z you will find z=0 , x=-1 ,y=6